OpenStudy (anonymous):

factor please help 1.a^3-125= 2.2x^2+X-4= 3.3X^2+X-4= 4.X^4negative2X^2-3=

OpenStudy (mathmale):

I attempted to help you with this very same post several hours ago, but you never responded. Why close that post and start a new one now?

OpenStudy (mathmale):

a^3-125 = a^3 - 5^3, which is a difference of two cubes. What's the formula for factoring the difference of two cubes?

OpenStudy (anonymous):

I do not no the formula @ mathmale

OpenStudy (mathmale):

the formula for factoring the difference of two squares is x^2 - a^2 = (x-a)(x+a). Find the equivalent formula for factoring the difference of two cubes, please.

OpenStudy (jdoe0001):

mathmale easy answer on that many do that, since they're just "waiting" to get a dry answer they can transcribe over =)

OpenStudy (mathmale):

Please do a Google search right now for "factoring the difference of two cubes."

OpenStudy (anonymous):

I'm doing the google search

OpenStudy (mathmale):

Great. Then you'll be able to finish this formula: a^3 - b^3 = ( - )( + - )

OpenStudy (anonymous):

(A+B)(A^2-AB+B^2)